# Self-Supervised Learning in Electron Microscopy: Towards a Foundation Model for Advanced Image Analysis

### v1

@@ Abstract
In this work, we explore the potential of self-supervised learning from unlabeled electron microscopy datasets, taking a step toward building a foundation model in this field. We show how self-supervised pretraining facilitates efficient fine-tuning for a spectrum of downstream tasks, including semantic segmentation, denoising, noise & background removal, and super-resolution. Experimentation with varying model complexities and receptive field sizes reveals the remarkable phenomenon that fine-tuned models of lower complexity consistently outperform more complex models with random weight initialization. We demonstrate the versatility of self-supervised pretraining across various downstream tasks in the context of electron microscopy, allowing faster convergence and better performance. We conclude that self-supervised pretraining serves as a powerful catalyst, being especially advantageous when limited annotated data are available and efficient scaling of computational cost are important.
@@ 1 Introduction
Microscopy, a fundamental tool in scientific research for several centuries, encompasses various branches, including optical, electron, scanning probe, and X-ray microscopy (Leng, 2013). Deep learning has become a universal tool for various downstream tasks in microscopy. So far, each specific imaging task is typically solved by a highly specialized model due to the absence of suitable foundation models (Subramanian et al., 2024). Foundation models additionally have the advantage that comprehensive and specialized training datasets are not required, which, however, comes at the cost of requiring larger training datasets. In this paper, our goal is to address and discuss these challenges. As a demonstrator, we chose a range of imaging problems from the field of electron microscopy. Our results, though, are applicable in more generality.
@@
Electron microscopy (EM) is a technique that uses a beam of accelerated electrons to obtain high-resolution images of biological as well as non-biological specimens. Applications of this technique exist across various scientific domains, including biology, materials science, nanotechnology, and physics. In the field of biology, EM has been used for studying a wide range of biological samples such as lungs, muscles, bones, or nerve tissue (Ke, 1971). In the realm of materials science, it has been utilized for visualization of the growth and characterization of nano- and microstructures (Lupan et al., 2014), orientation mapping of semicrystalline polymers (Panova et al., 2016), and the identification of crystal lattice defects (Oh et al., 2009).
@@
Imaging techniques and statistical analysis in EM have been instrumental in providing insights into the structure and properties of materials at various scales. Statistical analysis and classical machine learning methods have been used to analyze nanoparticles (Ezzat & Bedewy, 2020; Lee et al., 2020), identify defects in metals (Steinberger et al., 2023; Zhang et al., 2022a), and enhancing the quality of superresolution results in correlative tomography (Salas et al., 2017). These methods, however, have shortcomings, such as limited resolution, time-consuming sample preparation, and the need for expert interpretation of results.
@@
Deep learning (DL) and computer vision have been increasingly employed to address these limitations, enhance the capabilities of EM, and overcome the limitations of classical imaging and analysis methods by providing automated analysis, improved resolution, and enhanced interpretation of complex data: DL enables the extraction of valuable information from large datasets and offers new opportunities for quantitative image analysis in EM (Angermueller et al., 2016). It has been used for analyzing nanoparticles in TEM images (Sytwu et al., 2022), denoising TEM images (Vincent et al., 2021), identifying clean graphene areas (Sadre et al., 2021), automatically segmenting and tracking of crystalline defects, (Ruzaeva et al., 2023; Govind et al., 2024), decoding crystallography from high-resolution electron imaging and diffraction datasets (Aguiar et al., 2019), understanding important features of DL models for segmentation of high-resolution transmission electron microscopy (TEM) images (Horwath et al., 2020), and segmentation in large-scale cellular EM (Aswath et al., 2023), focused ion-beam scanning EM (FIB-SEM) (Khadangi et al., 2021) and high-resolution TEM data (Groschner et al., 2021).
@@
Conventional DL methods, such as convolutional neural networks, require large annotated datasets to be able to learn and generalize well on unseen examples. Manual annotation of datasets, especially EM images, is a time-consuming and labor-intensive task. To alleviate this, techniques such as transfer learning and self-supervised learning can be used. These techniques offer significant advantages for a range of computer vision tasks by enabling models to obtain general features from extensive datasets, facilitating knowledge transfer to specific tasks with limited labeled data. These approaches significantly reduce annotation costs, mitigate the problem of data scarcity, and strongly enhance generalization to unseen scenarios. Pretrained models exhibit faster convergence during fine-tuning, possess broader applicability across tasks, and provide resource-efficient solutions. The robust representations acquired through self-supervised learning contribute to improved performance in real-world scenarios, establishing it as an essential strategy in computer vision tasks across diverse domains (He et al., 2020; Valverde et al., 2021; Srinivas et al., 2022; Paul et al., 2016).
@@
Self-supervised learning aims to learn representations from the data itself without explicit manual supervision. It can be utilized to pretrain a model on a large amount of unlabeled data, allowing it to learn general features and representations from the data. These learned representations can then be transferred and fine-tuned for a specific task, effectively leveraging the knowledge gained from the self-supervised pretraining to improve performance on the target task. pretraining models on these tasks with unlabeled data and using the pretrained weights to fine-tune models on supervised tasks with limited annotations help improve model performance and reach faster convergence (Pathak et al., 2016; Chen et al., 2020b; He et al., 2020). The first step in self-supervised learning (pretraining on unlabeled data) is called the pretext. The second step (fine-tuning the pretrained models on annotated data) is called downstream.
@@ 2 Related Work
A commonly used approach in deep learning is pretraining models on large labeled datasets and fine-tuning on smaller datasets with limited annotations (Ridnik et al., 2021; Zhang et al., 2020; Marmanis et al., 2015). Such an approach usually performs well when the source data for pretraining is from a domain similar to the one from which the target data was obtained. In domains where large labeled datasets are scarce but an abundance of unlabeled datasets is available, self-supervised learning proves to be effective. Self-supervised learning leverages unlabeled datasets for pretraining, and the learned knowledge is then transferred to supervised downstream tasks with labeled data. Examples of successful self-supervised learning methods include contrastive learning, jigsaw puzzles, autoencoders, masked image modeling, and generative-based methods. SimCLR (Chen et al., 2020b) is a prominent self-supervised learning method that has demonstrated significant advancements in self-supervised learning on large-scale benchmarks such as ImageNet. It leverages a contrastive pretraining objective, which involves maximizing agreement between differently augmented views of the same data point while minimizing agreement with views from other data points. This approach has been shown to learn semantically meaningful representations from unlabeled data, making it a powerful method for self-supervised learning. Momentum Contrast (MoCo) (He et al., 2020) is another well-known technique that leverages contrastive learning for unsupervised visual representation learning. It has been widely applied in various domains, including remote sensing scene classification (Alosaimi et al., 2023), chest X-ray model pretraining (Sowrirajan et al., 2020), hand shape estimation (Zimmermann et al., 2021), and speaker embedding (Ding et al., 2020). The method has also been compared with other self-supervised learning techniques, demonstrating its effectiveness in learning representations from images and its potential for various downstream tasks (Zhang et al., 2022b). Conrad & Narayan used this method to pretrain DL models on cellular EM images.
@@
Masked image modeling, another self-supervised learning method, involves training a model to predict the original content of an image from a corrupted or masked version (Xie et al., 2022). This approach has been applied in various domains, including medical imaging and spectroscopic data identification. Li et al. proposed RGMIM (Region-Guided Masked Image Modeling) for COVID-19 detection, showcasing the potential of masked image modeling in medical imaging applications. Furthermore, Xue et al. highlighted the success of masked image modeling in self-supervised learning, demonstrating its ability to alleviate data-hungry issues and achieve competitive results. These examples underscore the significance of masked image modeling in learning robust representations and its applicability across diverse domains.
@@
Using GANs for self-supervised pretraining is also very effective. Chen et al. uses a GAN-based model to pretrain a model that learns image rotation. Guo et al. uses GAN-based pretraining for learning image similarity in remote sensing images. Other notable research in this area includes latent transformation detection (Patel et al., 2021), GAN-based image colorization for self-supervised visual feature learning (Treneska et al., 2022), and self-supervised learning for semantic segmentation of archaeological monuments (Kazimi & Sester, 2023).
In recent years, the prominence of foundation models in DL has surged. Foundation models are large, pretrained neural networks that are designed to perform a wide range of tasks across various domains. They are trained on massive datasets using DL techniques, such as self-supervised learning, to create a general-purpose model that can be adapted to a diverse set of distinctive tasks. They have proven to be effective in natural language processing (NLP) (Radford & Narasimhan, 2018), computer vision (Chen et al., 2020a), and other fields, providing a basis for the development of more specialized and task-specific models. A well-known example is the Segment Anything Model (SAM) (Kirillov et al., 2023), which is a groundbreaking work for image segmentation. It demonstrates impressive zero-shot performance on various tasks, often surpassing prior fully supervised results. Subramanian et al. investigates the utility of foundation models for scientific machine learning applications. The research reveals that appropriately fine-tuned pretrained models significantly reduce the need for annotated data in downstream tasks. Another step towards foundation models is taken by Liu et al., which uses self-supervised learning to mitigate the problem of class imbalance in datasets.
@@
Recently, pretraining methodologies have been explored in the domain of EM. In particular, the microstructure segmentation with DL encoders pretrained on a large microscopy dataset (Stuckner et al., 2022), classification of scanning electron microscope images of pharmaceutical excipients using deep convolutional neural networks with transfer learning (Iwata et al., 2022) and the unsupervised pretraining, the Momentum Contrast (MoCoV2) algorithm (Chen et al., 2020c) was used by Conrad & Narayan.
@@
In this paper, we take a first step towards foundation models for EM datasets and explore the application of self-supervised pretraining based on GANs, specifically the Pix2Pix architecture (Isola et al., 2017). The GAN model is pretrained on a large unlabeled Cellular Electron Microscopy (CEM) dataset called CEM500K (Conrad & Narayan, 2021). The pretrained generator model can be fine-tuned on a wide range of downstream tasks in EM, including semantic segmentation of nanoparticles in TEMs, denoising, noise & background removal, and super-resolution. We show that such a pretraining approach leads to faster convergence and higher predictive power on all of the mentioned tasks with limited annotated datasets. We also find that fine-tuning with pretrained weights helps smaller and architecturally simpler models achieve similar or even higher scores compared to training with random weight initialization.
@@ 3.1.1 CEM500K
CEM500K (Conrad & Narayan, 2021) is a large-scale, heterogeneous, unlabeled cellular EM image dataset developed for DL applications. The dataset is curated from experiments and various publicly available sources, encompassing 2D and 3D cellular EM images with diverse imaging modalities, sample preparation protocols, resolutions, and cell types. It includes examples from reconstructed FIB-SEM volumes, transmission EM (TEM) images, and EM image volumes and 2D images from various sources.
@@ 3.1.2 HRTEM Au Dataset
The Gold nanoparticle image dataset consists of high- and low-resolution TEM images of Gold (Au) nanoparticles with varying sizes (, , , and ) and different surface ligands, i.e., citrate (for ) and tannic acid. The images were acquired using an aberration-corrected TEAM 0.5 TEM for 2.2, 5, and nanoparticles, while low-resolution images of nanoparticles were obtained with a non-aberration-corrected TitanX TEM. The nanoparticles have a different crystalline structure than the embedding matrix (which is amorphous). This is the reason why the atomic arrangements look different in TEM. An important task in materials science is to segment such nanoparticles. The dataset was manually segmented and labeled, followed by preprocessing steps such as outlier removal and image standardization. To optimize memory usage during training, images were divided into 512 × 512-pixel patches, excluding patches consisting solely of amorphous background to address potential class imbalance issues (Sytwu et al., 2022). In this paper, we will refer to the datasets as “Au2.2nm”, “Au5nm”, “Au10nm”,“Au20nm”. Additionally, the dataset of Au nanoparticles (Groschner et al., 2021) was included in our experiments and is referred to as “Au5nmV1”. Examples of low and high-resolution TEM images of Gold nanoparticles and the corresponding ground truth segmentation annotations are shown in Figure 2.
@@ 3.1.3 TemImageNET
TEMImageNet is an open-source atomic-scale scanning transmission electron microscopy (ADF-STEM) image dataset. The dataset includes ten types of ground truth labels for training and validating DL models for tasks such as segmentation, super-resolution, background subtraction, denoising, and localization. The dataset comprises simulated ADF-STEM images of eight materials projected along multiple orientations with diverse atomic structures and crystallographic orientations. To replicate real-world experimental conditions, the images are augmented with realistic scan and Poisson noise, along with randomized linear and nonlinear low-frequency background patterns (Lin et al., 2021). The example simulated image and the ground truth labels are shown in Figure 3. The number of images in each dataset for training, validation and testing is given in Table 1.
@@ 3.2 Pretraining Method
We employ a GAN-based approach for pretraining on unlabeled data. GANs are originally designed for generating new data samples that resemble a given dataset. The GAN architecture involves two neural networks: a generator and a discriminator. These are trained simultaneously through adversarial training. The generator takes random noise as input and generates synthetic data samples that are indistinguishable from real data. The discriminator evaluates the real and generated data and distinguishes them from each other. Both networks are simultaneously trained in an adversarial fashion: the generator tries to improve its ability to generate realistic data to fool the discriminator. The discriminator, in turn, strives to become better at distinguishing between real and generated samples (Goodfellow et al., 2014).
Conditional Generative Adversarial Networks (cGANs) are an extension of the traditional GAN framework, where the generator is conditioned on additional information, typically in the form of class labels or other auxiliary data. The key idea is to guide the generation process based on specific conditions, allowing a more controlled and targeted generation of samples. In a cGAN, both the generator and the discriminator receive additional input information (conditioning) alongside the random noise for the generator and real/fake labels for the discriminator. The conditioning information could be anything relevant to the desired output, e.g., class labels, attributes, or other types of data. Conditional GANs have been used for various tasks, including image-to-image translation, image synthesis with specific attributes, and generating samples from certain classes. They provide a way to control and manipulate the characteristics of the generated data by incorporating additional information during the training process (Mirza & Osindero, 2014). The same loss function for cGANs as mentioned in (Isola et al., 2017) and also shown in Equation 3 in the Appendix is used.
In this paper, we use the cGAN model called Pix2Pix (Isola et al., 2017) for pretraining on the unlabeled CEM500K dataset of EM images. The images are fed to the generator with added noise, and the goal is to generate output images that are indistinguishable from the original ones. The trained generator model can then be fine-tuned on supervised downstream tasks explained in Section 3.3. To investigate the influence of the size of unlabeled datasets used for pretraining on the results of fine-tuning in the downstream tasks, we run pretraining experiments with different numbers of examples from the CEM500K dataset: 50K, 100K, and 200K examples.

### v2

# Self-Supervised Learning with Generative Adversarial Networks for Electron Microscopy

@@ Abstract
In this work, we explore the potential of self-supervised learning with Generative Adversarial Networks (GANs) for electron microscopy datasets. We show how self-supervised pretraining facilitates efficient fine-tuning for a spectrum of downstream tasks, including semantic segmentation, denoising, noise & background removal, and super-resolution. Experimentation with varying model complexities and receptive field sizes reveals the remarkable phenomenon that fine-tuned models of lower complexity consistently outperform more complex models with random weight initialization. We demonstrate the versatility of self-supervised pretraining across various downstream tasks in the context of electron microscopy, allowing faster convergence and better performance. We conclude that self-supervised pretraining serves as a powerful catalyst, being especially advantageous when limited annotated data are available and efficient scaling of computational cost is important.
@@ 1 Introduction
Microscopy, a fundamental tool in scientific research for several centuries, encompasses various branches, including optical, electron, scanning probe, and X-ray microscopy [24]. Electron microscopy (EM) is a technique that uses a beam of accelerated electrons to obtain high-resolution images of biological as well as non-biological specimens. Applications of this technique exist across various scientific domains, including biology, materials science, nanotechnology, and physics. In the field of biology, EM has been used for studying a wide range of biological samples such as lungs, muscles, bones, or nerve tissue [21]. In materials science, it has been utilized for visualization of the growth and characterization of nano- and microstructures [27], orientation mapping of semicrystalline polymers [32], and the identification of crystal lattice defects [31].
@@
Imaging techniques and statistical analysis in EM have been instrumental in providing insights into the structure and properties of materials at various scales. Statistical analysis and classical machine learning methods have been used to analyze nanoparticles [11, 23], identify defects in metals [43, 53], and enhancing the quality of superresolution results in correlative tomography [40]. These methods, however, have shortcomings, such as limited resolution, time-consuming sample preparation, and the need for expert interpretation of results.
@@
Deep learning (DL) and computer vision have been increasingly employed to address these limitations, enhance the capabilities of EM, and overcome the limitations of classical imaging and analysis methods by providing automated analysis, improved resolution, and enhanced interpretation of complex data: DL enables the extraction of valuable information from large datasets and offers new opportunities for quantitative image analysis in EM [3]. It has been used for analyzing nanoparticles in TEM images [46], denoising TEM images [49], identifying clean graphene areas [39], automatically segmenting and tracking of crystalline defects, [38, 13], decoding crystallography from high-resolution electron imaging and diffraction datasets [1], understanding important features of DL models for segmentation of high-resolution transmission electron microscopy (TEM) images [17], and segmentation in large-scale cellular EM [4], focused ion-beam scanning EM (FIB-SEM) [22] and high-resolution TEM data [14].
@@
Conventional DL methods, such as convolutional neural networks, require large annotated datasets to be able to learn and generalize well on unseen examples. Manual annotation of datasets, especially EM images, is a time-consuming and labor-intensive task. To alleviate this, techniques such as transfer learning and self-supervised learning can be used. These techniques offer significant advantages for a range of computer vision tasks by enabling models to obtain general features from extensive datasets, facilitating knowledge transfer to specific tasks with limited labeled data. These approaches significantly reduce annotation costs, mitigate the problem of data scarcity, and strongly enhance generalization to unseen scenarios. Pretrained models exhibit faster convergence during fine-tuning, possess broader applicability across tasks, and provide resource-efficient solutions. The robust representations acquired through self-supervised learning contribute to improved performance in real-world scenarios, establishing it as an essential strategy in computer vision tasks across diverse domains [16, 48, 42, 35].
@@
Self-supervised learning aims to learn representations from the data itself without explicit manual supervision. It can be utilized to pretrain a model on a large amount of unlabeled data, allowing it to learn general features and representations from the data. These learned representations can then be transferred and fine-tuned for a specific task, effectively leveraging the knowledge gained from the self-supervised pretraining to improve performance on the target task. Pretraining models on these tasks with unlabeled data and using the pretrained weights to fine-tune models on supervised tasks with limited annotations help improve model performance and reach faster convergence [34, 7, 16]. The first step in self-supervised learning (pretraining on unlabeled data) is called the pretext. The second step (fine-tuning the pretrained models on annotated data) is called downstream.
@@ 2 Related Work
A commonly used approach in deep learning is pretraining models on large labeled datasets and fine-tuning on smaller datasets with limited annotations [36, 54, 29]. Such an approach usually performs well when the source data for pretraining is from a domain similar to the one from which the target data was obtained. In domains where large labeled datasets are scarce but an abundance of unlabeled datasets is available, self-supervised learning proves to be effective. Self-supervised learning leverages unlabeled datasets for pretraining, and the learned knowledge is then transferred to supervised downstream tasks with labeled data. Examples of successful self-supervised learning methods include contrastive learning, jigsaw puzzles, autoencoders, masked image modeling, and generative-based methods. SimCLR [7] is a prominent self-supervised learning method that has demonstrated significant advancements in self-supervised learning on large-scale benchmarks such as ImageNet. It leverages a contrastive pretraining objective, which involves maximizing agreement between differently augmented views of the same data point while minimizing agreement with views from other data points. This approach has been shown to learn semantically meaningful representations from unlabeled data, making it a powerful method for self-supervised learning. Momentum Contrast (MoCo) [16] is another well-known technique that leverages contrastive learning for unsupervised visual representation learning. It has been widely applied in various domains, including remote sensing scene classification [2], chest X-ray model pretraining [41], hand shape estimation [56], and speaker embedding [10]. The method has also been compared with other self-supervised learning techniques, demonstrating its effectiveness in learning representations from images and its potential for various downstream tasks [55]. Conrad and Narayan used this method to pretrain DL models on cellular EM images.
@@
Masked image modeling, another self-supervised learning method, involves training a model to predict the original content of an image from a corrupted or masked version [51]. This approach has been applied in various domains, including medical imaging and spectroscopic data identification. Li et al. proposed RGMIM (Region-Guided Masked Image Modeling) for COVID-19 detection, showcasing the potential of masked image modeling in medical imaging applications. Furthermore, Xue et al. highlighted the success of masked image modeling in self-supervised learning, demonstrating its ability to alleviate data-hungry issues and achieve competitive results. Caron et al. show the prominence of self-supervised learning on various tasks using vision transforms. These examples underscore the significance of masked image modeling in learning robust representations and its applicability across diverse domains.
@@
Using GANs for self-supervised pretraining is also very effective. Chen et al. uses a GAN-based model to pretrain a model that learns image rotation. Guo et al. uses GAN-based pretraining for learning image similarity in remote sensing images. Other notable research in this area includes latent transformation detection [33], GAN-based image colorization for self-supervised visual feature learning [47], and self-supervised learning for semantic segmentation of archaeological monuments [20].
@@
Recently, pretraining methodologies have been explored in the domain of EM. In particular, the microstructure segmentation with DL encoders pretrained on a large microscopy dataset [44], classification of scanning electron microscope images of pharmaceutical excipients using deep convolutional neural networks with transfer learning [19] and the unsupervised pretraining, the Momentum Contrast (MoCoV2) algorithm [8] was used by Conrad and Narayan.
@@
In this paper, we explore the application of self-supervised pretraining based on GANs, specifically the Pix2Pix architecture [18], for EM images. The GAN model is pretrained on a large unlabeled Cellular Electron Microscopy (CEM) dataset called CEM500K [9]. The pretrained generator model can be fine-tuned on a wide range of downstream tasks in EM, including semantic segmentation of nanoparticles in TEMs, denoising, noise & background removal, and super-resolution. We show that such a pretraining approach leads to faster convergence and higher predictive power on all of the mentioned tasks with limited annotated datasets. We also find that fine-tuning with pretrained weights helps smaller and architecturally simpler models achieve similar or even higher scores compared to training with random weight initialization.
@@ 3.1.1 CEM500K
CEM500K [9] is a large-scale, heterogeneous, unlabeled cellular EM image dataset developed for DL applications. The dataset is curated from experiments and various publicly available sources, encompassing 2D and 3D cellular EM images with diverse imaging modalities, sample preparation protocols, resolutions, and cell types. It includes examples from reconstructed FIB-SEM volumes, transmission EM (TEM) images, and EM image volumes and 2D images from various sources.
@@ 3.1.2 HRTEM Au Dataset
The Gold nanoparticle image dataset consists of high- and low-resolution TEM images of Gold (Au) nanoparticles with varying sizes (, , , and ) and different surface ligands, i.e., citrate (for ) and tannic acid. The images were acquired using an aberration-corrected TEAM 0.5 TEM for 2.2, 5, and nanoparticles, while low-resolution images of nanoparticles were obtained with a non-aberration-corrected TitanX TEM. The nanoparticles have a different crystalline structure than the embedding matrix (which is amorphous). This is the reason why the atomic arrangements look different in TEM. An important task in materials science is to segment such nanoparticles. The dataset was manually segmented and labeled, followed by preprocessing steps such as outlier removal and image standardization. To optimize memory usage during training, images were divided into 512 × 512-pixel patches, excluding patches consisting solely of amorphous background to address potential class imbalance issues [46]. In this paper, we will refer to the datasets as “Au2.2nm”, “Au5nm”, “Au10nm”,“Au20nm”. Additionally, the dataset of Au nanoparticles [14] was included in our experiments and is referred to as “Au5nmV1”. Examples of low and high-resolution TEM images of Gold nanoparticles and the corresponding ground truth segmentation annotations are shown in Figure 2.
@@ 3.1.3 TemImageNET
TEMImageNet is an open-source atomic-scale scanning transmission electron microscopy (ADF-STEM) image dataset. The dataset includes ten types of ground truth labels for training and validating DL models for tasks such as segmentation, super-resolution, background subtraction, denoising, and localization. The dataset comprises simulated ADF-STEM images of eight materials projected along multiple orientations with diverse atomic structures and crystallographic orientations. To replicate real-world experimental conditions, the images are augmented with realistic scan and Poisson noise, along with randomized linear and nonlinear low-frequency background patterns [26]. The example simulated image and the ground truth labels are shown in Figure 3. The number of images in each dataset for training, validation and testing is given in Table 1.
@@ 3.2 Pretraining Method
We employ a GAN-based approach for pretraining on unlabeled data. GANs are originally designed for generating new data samples that resemble a given dataset. The GAN architecture involves two neural networks: a generator and a discriminator. These are trained simultaneously through adversarial training. The generator takes random noise as input and generates synthetic data samples that are indistinguishable from real data. The discriminator evaluates the real and generated data and distinguishes them from each other. Both networks are simultaneously trained in an adversarial fashion: the generator tries to improve its ability to generate realistic data to fool the discriminator. The discriminator, in turn, strives to become better at distinguishing between real and generated samples [12].
Conditional Generative Adversarial Networks (cGANs) are an extension of the traditional GAN framework, where the generator is conditioned on additional information, typically in the form of class labels or other auxiliary data. The key idea is to guide the generation process based on specific conditions, allowing a more controlled and targeted generation of samples. In a cGAN, both the generator and the discriminator receive additional input information (conditioning) alongside the random noise for the generator and real/fake labels for the discriminator. The conditioning information could be anything relevant to the desired output, e.g., class labels, attributes, or other types of data. Conditional GANs have been used for various tasks, including image-to-image translation, image synthesis with specific attributes, and generating samples from certain classes. They provide a way to control and manipulate the characteristics of the generated data by incorporating additional information during the training process [30]. The same loss function for cGANs as mentioned in [18] is used.
In this paper, we use the cGAN model called Pix2Pix [18] for pretraining on the unlabeled CEM500K dataset of EM images. The images are fed to the generator with added noise, and the goal is to generate output images that are indistinguishable from the original ones. The trained generator model can then be fine-tuned on supervised downstream tasks explained in Section 3.3. To investigate the influence of the size of unlabeled datasets used for pretraining on the results of fine-tuning in the downstream tasks, we run pretraining experiments with different numbers of examples from the CEM500K dataset: 50K, 100K, and 200K examples.